Circular No. 7082

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1998fc AND 1998fd IN ANONYMOUS GALAXIES A. Gal-Yam and D. Maoz, Tel Aviv University, report for the Wise Ob-

servatory Optical Transients Search (cf. IAUC 7055) their discovery of two apparent supernovae on unfiltered CCD images taken with the Wise Observatory 1-m telescope on 1998 Dec. 20 and 24. SN 1998fc, discovered at R = 20.5 on Dec. 20, is located at α = 2^{h}59^{m}12^{s}.61, δ = +3^{o}29′39′′.0 (equinox 2000.0), which is 21′′ west and 16.7′′ south of the nucleus of the giant galaxy at the center of the galaxy cluster Abell 403 (z = 0.103). Summation of several previous images of this field does not reveal any galaxy that is a more plausible host for SN 1998fc. SN 1998fc was invisible in images obtained on 1998 Nov. 14 (limiting mag R ∼ 22.5) but was confirmed (R ∼ 21.2) on CCD images taken by J. Dan and R. Beeri on 1999 Jan. 7 and 9 with the same telescope. SN 1998fd, discovered at R = 21.3 on 1998 Dec. 24, is located at α = 1^{h}19^{m}18^{s}.06, δ = +15^{o}55′24′′.1, which is 5′′.8 east and 2′′.8 north of the host galaxy center. SN 1998fd was also confirmed on images taken on 1999 Jan. 9 by Beeri and shows some fading. SN 1998fd was invisible in images obtained on 1998 Nov. 15 (limiting mag R ∼ 22.5). Finding charts can be

obtained via ftp://wise-obs.tau.ac.il/pub/avishay/sn33.fits (SN

1998fc) and /sn32.fits (SN 1998fd).

SUPERNOVA 1998eu IN ANONYMOUS GALAXY Gal-Yam and Maoz also write: “A spectrum of SN 1998eu (IAUC 7055)

was obtained by R. A. Stathakis, R. D. Cannon, and D. M. James, Anglo-Australian Observatory, on 1998 Dec. 11. The spectrum (range 350–625 nm), obtained with the 4-m Anglo-Australian Telescope (+ RGO spectrograph and 300B grating), shows narrow emission lines of [O II] 372.7-nm and [O III] 500.7 nm, setting the galaxy redshift at z = 0.181, and consistent with it being a member of the galaxy cluster Abell 125 (z = 1.88). Four broad (FWHM 10.0-nm) emission peaks at rest wavelengths of 395.0, 427.0, 459.5, and 493.0 nm (assuming the above redshift) are detected. The spectrum is similar to that of a type-Ia supernova that is ∼ 4 weeks old, consistent with the discovery date of SN 1998eu (Nov. 14).”

RX J0052.1–7319 P. Kahabka, University of Amsterdam, reports an improved position

for the 15.3-s x-ray pulsar in the Small Magellanic Cloud (cf. IAUC 7081), derived from a ROSAT HRI observation on 1995 May 18–June 3: α = 0^{h}52^{m}15^{s}.5, δ = −73^{o}19′14′′ (equinox 2000.0; ± 4′′ at 90 percent confidence).

1999 January 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green